OpenStudy (anonymous):

−3|2x + 6| = −12 I just don't understand these questions

OpenStudy (anonymous):

What don't you understand?

OpenStudy (anonymous):

I am unsure how to solve it. I tried already and got most of the questions wrong.

OpenStudy (anonymous):

I know the answer but I don't know the exact formula, it's weird with me how I just know the answers but I don't solve them like everybody else does. I'll look up a formula and go over it.

OpenStudy (anonymous):

I do that a lot too. The main part I don't understand is how there can be two answers.

OpenStudy (anonymous):

There is a negative and a positive answer because everything in an absolute value sign has to come out to be equal, so let's say x is -1, we would do -1x2 = -2+6 = 4. and 4x-3 = -12

OpenStudy (mathstudent55):

The first step is to divide both sides by -3. Try that and show what you get.

OpenStudy (anonymous):

Yea listen to this guy, I'm not good at explaining, just good at getting the answers lol

OpenStudy (mathstudent55):

@MiaS The first step is to divide both sides by -3. Try that and show what you get.

OpenStudy (anonymous):

For the 12 i got 4 On the other side how do i divide that? Just remove the negative 4?

OpenStudy (mathstudent55):

−3|2x + 6| = −12 |2x + 6| = 4

OpenStudy (mathstudent55):